Find the limit of Xn=1/2[Xn-1+Xn]





1
Expert's answer
2022-03-19T02:41:10-0400

"xn=\\frac{1}{2}(X\ufeff\ufeffn-1 +xn)"

"\lim xn=x1+lim \sum_{k=1}^n(x_{k+1}-x_k)"

"=x_1+x_2-x_1+lim \sum_{k=2}^n\frac{(-1)^{k-1}c}{2^{k-1}}"





"=x_2+ lim \sum_{k=1}^{\frac{n}{2}}\frac{c}{2^{2k}}-lim \sum_{k=1}^{\frac{n}{2}}\frac{c}{2^{2k-1}}"

"=x_2+c lim(\sum_{k=1}^{\frac{n}{2}}\frac{1}{4^k}-\sum_{k=1}^{\frac{n}{2}}\frac{1}{2}\frac{1}{2^{2k-1}})"

"=x_2+c(\frac{1}{4}- (\sum_{k=0}^{\frac{n-1}{2}}\frac{1}{2}\frac{1}{4^k})"





"=x_2+c(\frac{1}{4}-\frac{2}{3})
=x_2+\frac{2}{3}c"